OpenStudy (thewalkingdeadrick):

A recipe that makes 7 servings calls for 1 1/6 cups of juice. How many cups of juice are there per serving?

OpenStudy (thewalkingdeadrick):

@Nnesha

OpenStudy (thewalkingdeadrick):

@mathmath333

OpenStudy (thewalkingdeadrick):

WILL MEDAL AND FAN!!!

OpenStudy (thewalkingdeadrick):

@THEHELPER123

OpenStudy (anonymous):

1/6

OpenStudy (thewalkingdeadrick):

thank you @JasperRayWolf-Alysa88

OpenStudy (anonymous):

no problem

OpenStudy (texaschic101):

1 1/6 = 7/6 7/(7/6) = 1/x cross multiply 7x = 7/6 -- divide both sides by 7 x = (7/6)/7 x = 7/6 * 1/7 x = 7/42 which reduces to 1/6